# Explainable AI for Autism Diagnosis: Identifying Critical Brain Regions Using fMRI Data

### v1

@@ III-B Feature Extraction
To reveal the patterns of coordination between brain regions, the data is converted to a functional connectivity matrix (FCM) using Pearson’s correlation coefficient $R(i,j)$ between the intensity time series from two regions of interests (ROIs). Furthermore, Fisher’s Z transformation  [^4] Fisher’s Z transformation is a statistical method used to convert Pearson’s correlation coefficient to a normally distributed variable. This transformation is particularly useful when working with correlation coefficients, as it stabilises their variance and makes them more suitable for statistical analyses. was applied to the data.
The Fisher’s z-transformed correlation coefficients were then represented as a 116 x 116 matrix which was symmetric with respect to the diagonal due to the symmetric nature of correlation matrices $R(i,j)=R(j,i)$. To reduce redundancy, the lower triangular matrix of the FCM was taken and flattened to form a feature vector for each of the $884$ samples creating a feature vector with $6670$ features.
@@
Support Vector Machine with Recursive Feature Elimination (SVM-RFE) was used to reduce the feature set from $6670$ features to a 1000 features per sample [15]. This choice considers computational constraints, proven performance and the fact that SVM-RFE is a supervised feature selection approach, whereas unsupervised methods like Principal Component Analysis (PCA), operate without knowledge of the classification task, which can affect the accuracy.
@@ III-C Deep Learning Model
This study utilises a DL architecture (Figure 1) that comprises of a Stacked Sparse Autoencoder (SSAE) combined with a softmax classifier to classify ASD using functional connectivity data.
The SSAE was chosen for this study due to its ability to effectively handle high-dimensional functional connectivity data. Autoencoders are neural networks that learn to compress input data into a lower-dimensional representation and then reconstruct it, making them ideal for capturing essential features in complex datasets. The “sparse” aspect is especially important for ASD classification, as it encourages the model to focus on the most distinctive biomarkers within the functional connectivity data while reducing noise and overfitting. The hierarchial structure of the SSAE, consisting of multiple layers of sparse autoencoders, encourages the model to learn increasingly abstract features of the data with each layer. This is important since biomarkers in functional connectivity may involve complex multi-level patterns.
To address potential overfitting and improve generalisability, we incorporate a sparsity constraint using the Kullback-Leibler (KL) divergence  [^5] Kullback-Leibler (KL) divergence is a measure from information theory that quantifies how different one probability distribution is from another term in the loss function:
@@
where $MSE$ is the mean squared error between input $x$ and its reconstruction $\hat{x}$, $\rho$ is the desired sparsity, $\hat{\rho}$ is the average activation of a neuron, and $\beta$ controls the weight of the sparsity penalty.
Figure 1: The overall architecture of the proposed approach, incorporating both the SSAEs and softmax classifier.
Our SSAE consisted of two autoencoder layers that used the ReLU activation function. The first reduces 1000 input features to 500, and the second further compresses to 100 neurons. The output was then fed into a softmax classifier that was trained using cross-entropy loss for ASD/TC classification. This is visualised in Figure 1.
@@ III-D Training and Evaluation
The model was trained in two stages combining the strengths of both supervised learning and unsupervised learning. To ensure robust evaluation and training a stratified k-fold cross-validation approach with $k$ value of 5 was employed. A k-fold value of 5 was chosen to balance computational cost with variance reduction in performance estimates with regard to the dataset size of 884. A higher $k$ value led to there not being enough test cases for the model to be tested properly and a lower $k$ value leads to there being not enough training data and the model struggles to generalise. Within each fold, the dataset was split into training, validation, and testing sets, with careful consideration given to maintaining a balance between computational efficiency and robust performance assessment.
@@ III-D1 Unsupervised Pre-training
The SSAE underwent unsupervised pre-training using the greedy layer-wise training approach [57]. This sequential approach allows each autoencoder layer to learn increasingly abstract representations of the input data. SAE 1 was trained first and its compressed hidden layer representation then served as input for training SAE 2. Both layers were trained using the Adam optimiser, chosen for its adaptive learning rate and faster convergence [58]. Performance monitoring during pre-training combined Mean Squared Error (MSE) and a sparsity penalty achieved through KL Divergence. This approach has shown efficacy in ASD classification, potentially improving generalisation and reducing overfitting [56, 59, 15].
@@ III-D2 Supervised Fine-Tuning
Following pre-training, the SSAE’s encoder layers were combined with a softmax classifier for supervised fine-tuning. This stage optimises classification performance by backpropagating the error signal [^6] The error signal is the cross entropy loss from the softmax classifier, not the reconstruction loss used to train the encoder layers in the pre-training step. from the softmax classifier, modifying the encoder weights to improve discrimination between ASD and TC classes. Unlike many previous approaches that focused solely on fine-tuning the classifier, our study fine-tunes both the encoder layers and the classifier, which may contribute to higher classification accuracy [55, 56, 59].
The fine-tuning stage leveraged a hybrid training approach, combining the SSAE’s unsupervised representation learning capabilities with supervised training for the specific ASD classification task. This process employed the Adam optimiser [^7] Hyperparameters: Pre-training - 50 epochs, learning rate of 0.001, weight decay of 0.0001, batch size of 128, KL Divergence parameters: $\rho$ = 0.2, $\beta$ = 2. Fine-tuning - 50 epochs, learning rate of 0.0001, weight decay of 0.0001, batch size of 128., chosen for its demonstrated efficiency in the pre-training phase. The hyperparameters were carefully selected to balance performance gains against potential overfitting, allowing for subtle adjustments to the pre-trained weights while maintaining regularisation.
@@ III-D3 Performance Evaluation
A rigorous evaluation of the models’s performance on the ASD classification task is necessary to assess its effectiveness and generalisability. We used the previously established stratified k-folds and splits to maintain consistency. Key metrics, including accuracy, sensitivity, specificity, precision, and F1-score, were calculated for each fold and then averaged to provide an overall assessment of the model’s classification ability. Additionally, loss graphs were visualised for training and validation to avoid and monitor overfitting. This evaluation seeks to determine whether the model can surpass the performance limitations often observed in similar ASD classification models while avoiding overfitting.
@@ III-E1 Interpretability Methods
To identify potential neurobiological markers of ASD and enhance the model’s explainability, a rigorous feature analysis was conducted. This investigation involved multiple interpretability techniques, including:
  • Integrated Gradients [38]: Calculates feature importance by examining how changes in input values impact the predictions
  • LIME [36]: Approximates complex model decisions locally through perturbations of the input data.
  • SHAP [35]: A game-theoretic approach attributing feature importance based on its contribution to the model’s output.
  • DeepLift [34]: Decomposes the output prediction by backpropagating the contributions of neurons against a reference point.
  • DeepLiftShap [60]: Combines DeepLift’s insights with SHAP values for enhanced explainability.
  • GradientShap [35]: Computes SHAP values using gradients, offering computational advantages.
  • GuidedBackprop [61]: Highlights input features relevant to the prediction by visualising positive gradients during backpropagation.
These methods were chosen as they use diverse approaches to enhance explainability. Gradient-based techniques (Integrated Gradients, DeepLift, DeepLiftSHAP, GradientSHAP) directly examine feature influence on predictions. The LIME approach explores model sensitivity to input perturbations, while the SHAP approach employs a game-theoretic approach for feature attribution. Methods like DeepLiftSHAP and GuidedBackprop combine strengths for greater robustness. As each method may yield varying feature rankings due to their differing focus, a benchmark is crucial for trustworthy interpretation.
@@ III-E2 Remove and Retrain (ROAR)
To systematically compare these methods, the Remove and Retrain (ROAR) benchmark was employed [18]. ROAR works by:
1. Identifying Important Features: Each interpretability method generates a ranking of input features, showing which features the method considers most important for making predictions.
2. Testing Importance: ROAR tests how crucial those top features truly are. It takes the top ’t’ percent of features (using different threshold values like 20%, 40%, etc.) and replaces them with meaningless values.
3. Retraining and Comparing: For each threshold, the model is retrained on the modified data, and its performance is measured. If an interpretability method is reliable, removing its top-ranked features should cause a steep decline in the model’s performance compared to removing features identified by other methods. Once these are computed they are displayed as a line graph to see which method had the steepest decline and that method is noted as the best performing method. Additionally, the results are compared to a baseline where features are removed randomly. This ensures that any observed performance drops are due to the interpretability method identifying truly important features, and not just the general effect of removing any features.
For each interpretability method the top features were ranked, then during the ROAR analysis the top $t$ percent of features were replaced by zero. Zero was chosen as the uninformative value because the input features are correlation coefficients and zero correlation indicates that no positive or negative relationship exists between the two variables. Hence, zero is uninformative in the context of functional connectivity. The threshold values of $t$ were chosen to focus on studying the importance of the top 10% of features in ASD classification while still monitoring the rest of the features  [^8] The threshold values of $t$ chosen for this research are: $[0.01,0.05,0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,0.95,0.99]$. The results are analysed in section IV.
@@ III-E3 Important Regions of Interest (ROIs)
Figure 2: Important ROIs (yellow) identified by the Integrated Gradients approach when using CPAC pipeline.
Figure 3: Important ROIs (yellow) identified by the Integrated Gradients approach when using CCS pipeline.
@@
Once ROAR established the reliability of some interpretability methods, the focus was moved to the brain regions (ROIs) represented by the features identified by the method to understand their role in ASD.
Each feature of the model is a correlation coefficient which means each feature is a connection between two ROIs. To calculate individual important ROIs, weights were calculated by multiplying the frequency of that ROI in the top connections and its highest weight. These individual weights were also graphed with a larger size and brighter color of node indicating more importance showcased in Figures 4, 4 and 4. Once the most important ROIs for each preprocessing pipeline were determined, each region was mapped to its Brodmann Area and the overlap between the different preprocessing pipelines were studied.
@@ IV Results and Discussion
The model achieved state-of-the-art performance on the ABIDE dataset, with a maximum accuracy of 98.2% across three different preprocessing pipelines, surpassing previous models for ASD classification (Table II). Evaluation metrics, including specificity, precision, and F1-score, were high with values of 0.98, 0.98 and 0.97 respectively. Analysis of the model’s decision-making process identified a set of discriminative functional connectivity features. These results are presented in detail, alongside a discussion of model performance, interpretability performance, feature analysis and the potential significance of the identified ROIs.
IV-A Model performance
Figure 5: Loss graphs for our model in the first K-Fold. AE1, AE2 and classifier loss is reconstruction loss during unsupervised pre-training. Model loss and accuracy is during supervised fine-tuning.
Model Accuracy
Model without FD Filtering 91%
Model with FD Filtering 98.2%
TABLE I: Model accuracy with and without framewise displacement filtering of samples $>$ 0.2mm FD.
The model demonstrated exceptional performance on the ASD classification task, particularly when samples with a FD value $>$ of 0.2mm were filtered. The 7% decrease in accuracy (Table I) along with the observation that the autoencoders could not accurately reconstruct the data when a sample has more than 0.2mm FD, suggests that the hypothesis regarding FD’s impact on data quality is indeed valid. This finding aligns with previous research on the subject [52] [51]. In particular, all the incorrect predictions during testing also had a high FD, suggesting that head movement during MRI has a significant effect on the precision of the diagnosis of ASD. This finding contrasts with previously reported results on the subject [53].
Using FD Filtering, an accuracy of 98.2% using the DPARSF pipeline (averaged across 5 K-Folds) was achieved by the model. These metrics surpass all previous state-of-the-art models (Table II) using the fMRI modality, reported as 93.59% accuracy in [15]
Notably, CCS, DPARSF and CPAC all performed within 1% of each other while NIAK trailed behind in every metric by an average of 4% (Figure 6). This difference in performance likely stems from the different approaches used for NIAK compared to CCS, CPAC, and DPARSF. In the NIAK pipeline, the functional data was written into template space prior to calculating the statistical derivatives, whereas for the others the statistical derivatives were calculated in native space and then the functional data was written into template space [50].
@@ IV-A Model performance
TABLE II: Accuracy comparison with papers that have a dataset $>$ 500 samples using the fMRI modality.
The loss graphs (Figure 5) reveal an interesting distinction. During the unsupervised pre-training stage, the loss graphs of AE 1 and AE 2 show similar losses during training and validation. Since this is reconstruction loss and the AEs can accurately reconstruct the data, we can infer that the dataset is mostly homogenous and the distinctions between samples are minuscule when not differentiating between ASD and TC which is expected since the difference between a brain with ASD and one without is subtle [65]. During fine-tuning, we use a cross-entropy loss which results in a more typical graph where the validation loss is slightly higher than the training loss, but both are decreasing. This pattern suggests that the model is learning continuously without overfitting, which is further supported by the accuracy graph.
The model’s success in classifying ASD likely stems from its hybrid architecture. The combination of unsupervised pre-training and supervised fine-tuning appears to have enabled the model to learn meaningful representations of complex connectivity patterns. This aligns with findings suggesting that DL models benefit from such multi-stage training approaches, especially in medical image analysis [66].
@@ IV-B1 ROAR Benchmark
The ROAR analysis identified which interpretability methods are reliable. It showed that Integrated Gradients, Deep fLift, DeepLiftShap, and GradientShap methods all had similar results and showed the steepest drop in accuracy when the initial features were removed, indicating that the gradient-based methods identified the most discriminative features required for ASD classification (Figure 7). The Integrated Gradients approach showed the steepest drop in accuracy and was therefore assumed the most reliable method. Notably, the LIME approach did not perform much better than the random baseline, indicating that feature perturbation-based techniques do not perform well in the functional connectivity field. This contrasts with results found in the original ROAR study [18] that showed that the Integrated Gradients approach performed similarly to the random baseline, further emphasising that the best-performing interpretability methods change depending on the model architecture and the data’s modality.
@@ IV-B2 Feature Analysis
Feature analysis using the Integrated Gradients approach revealed a set of consistent brain connectivity patterns highly discriminative for the classification of ASD across the different preprocessing pipelines which can be observed in Figures 4, 4 and 4.
Important brain areas (parcellated per the AAL atlas), identified by the interpretation method are shown in columns CCS, CPAC, and DPARSF in Table III. The regions are arranged in descending order of importance score, with their corresponding Brodmann Areas [67]. The most consistently important regions across pipelines are areas 17 (Primary visual cortex) and 18 (Secondary visual cortex), with areas 41 and 42 (Auditory cortex) also showing significance.
@@
The convergence of our machine learning-based findings with these different studies provides strong validation that our model has indeed captured genuine neurobiological markers of ASD, rather than dataset-specific patterns. These results also demonstrate its potential to corroborate and extend existing knowledge in autism research. The consistency of our findings across different preprocessing pipelines and their alignment with diverse research approaches (genetic, neuroanatomical, and functional) highlights the robustness of our method and its potential for advancing our understanding of the neurobiological basis of ASD. Figure 8 highlights these important regions, providing a visual representation of the key areas identified by our analysis.
@@
Figure 8: Shows the Brodmann’s area[74]. The different colors represent different brain functions. The common important brain regions from the table III are highlighted.

### v2

@@ III-B Feature Extraction
To reveal the patterns of coordination between brain regions, each fMRI time-series was converted into a functional connectivity matrix (FCM) by calculating Pearson’s correlation coefficient $R(i,j)$ between the intensity time series from every pair of ROIs. We then applied Fisher’s Z transformation to normalise these correlation coefficients. Finally, to reduce redundancy, we flattened the lower triangular of each symmetric FCM $(116\times 116)$, yielding $6670$ features per sample.
@@
To address this, we employed Support Vector Machine with Recursive Feature Elimination (SVM-RFE) for dimensionality reduction. We chose SVM-RFE (over unsupervised methods like PCA) for its task-specific feature selection and proven performance in similar studies [12], ultimately reducing the feature set from 6,670 to 1,000 features per sample. This mitigates the computational burden and helps the model focus on the most discriminative functional connections.
Figure 1: The overall architecture of the model, incorporating both the SSAE and softmax classifier.
@@ III-C Deep Learning Model
This study utilises a DL architecture (Figure 1) comprising a Stacked Sparse Autoencoder (SSAE) combined with a softmax classifier to classify ASD using functional connectivity data. Autoencoders learn a compressed representation of the input, and the sparse constraint encourages focus on only the most distinctive biomarkers, which is important for high-dimensional functional connectivity data.
Our choice of a two-layer Stacked Sparse Autoencoder (SSAE) with ReLU activation and KL divergence [^3] Kullback-Leibler (KL) divergence is a measure from information theory that quantifies how different one probability distribution is from another sparsity penalty is motivated by both practical and empirical considerations [55, 12]. We found that two hidden layers strike an optimal balance between expressive power and training complexity, especially given our sample size. Hyperparameter tuning indicated that deeper networks did not improve performance, so this design effectively balances model expressiveness with the risk of overfitting. ReLU activations have shown robust performance in similar high-dimensional neuroimaging tasks, avoiding saturation issues observed in sigmoid/tanh functions. Crucially, the KL divergence term enforces neuron-level sparsity, highlighting discriminative features while mitigating overfitting [54], as captured by the following loss function $L(x,\hat{x})$ between input $x$ and its reconstruction $\hat{x}$:
@@
where $MSE$ is the mean squared error between input $x$ and its reconstruction $\hat{x}$, $\rho$ is the desired sparsity, $\hat{\rho}_{j}$ is the average activation of the $j$-th hidden neuron, and $\beta$ controls the weight of the sparsity penalty.
Our SSAE consisted of two autoencoder layers that used the ReLU activation function. The first reduces 1000 input features to 500, and the second further compresses to 100 neurons. The output was then fed into a softmax classifier that was trained using cross-entropy loss for ASD/TC classification. This is visualised in Figure (1).
@@ III-D Training and Evaluation
The model was trained in two stages—unsupervised pre-training followed by supervised fine-tuning—to combine the strengths of both approaches. We used stratified 5-fold cross-validation to balance computational cost with variance reduction, given our dataset of 884 samples. Within each fold, the data was split into 64% training, 16% validation, and 20% testing, ensuring distinct subsets at every stage. A public repository with our code will be provided to facilitate reproducibility.
@@ III-D1 Unsupervised Pre-training
The SSAE underwent unsupervised pre-training using the greedy layer-wise training approach [56]. This sequential approach allows each autoencoder layer to learn increasingly abstract representations of the input data. SAE 1 was trained first and its compressed hidden layer representation then served as input for training SAE 2. Each layer used Mean Squared Error (MSE) combined with a sparsity penalty via KL Divergence for reconstruction fidelity [55, 57, 12]. We employed the Adam optimiser [58] throughout, chosen for its adaptive learning rate and rapid convergence.
@@ III-D2 Supervised Fine-Tuning
Once pretraining was complete, the encoder layers were combined with a softmax classifier. In contrast to approaches that fine-tune only the classifier, we fine-tuned both the encoder and classifier parameters together to enhance discrimination between ASD and TC [54, 55, 57]. This supervised phase again used Adam, with hyperparameters [^4] Hyperparameters: Pre-training - 50 epochs, learning rate of 0.001, weight decay of 0.0001, batch size of 128, KL Divergence parameters: $\rho$ = 0.2, $\beta$ = 2. Fine-tuning - 50 epochs, learning rate of 0.0001, weight decay of 0.0001, batch size of 128. adjusted to avoid overfitting while refining the learned representations for classification.
@@ III-D3 Performance Evaluation
We report accuracy, sensitivity, specificity, precision, and F1-scores, each averaged across the five k-folds. This ensures a thorough assessment of model performance beyond simple accuracy. Additionally, we monitored training and validation losses to detect overfitting, providing insight into whether the model surmounts typical performance limitations reported in prior ASD classification work.
@@ III-E1 Interpretability Methods
This section explores the interpretability methods employed to reveal how the model arrives at its decisions. Interpretable features are crucial for clinical acceptance, as they help clarify the neural mechanisms underlying ASD classification and thus support potential clinical translation.
We evaluated seven approaches (gradient- and perturbation-based) to capture different perspectives of feature importance. Specifically, we used Integrated Gradients [39], LIME [37], SHAP [36], DeepLift [35], DeepLiftShap [59], GradientShap [36], and GuidedBackprop [60]. Gradient-based methods (Integrated Gradients, DeepLift, DeepLiftShap, GradientShap) directly quantify how changes in input features affect predictions. LIME employs local perturbations to approximate model decisions, whereas SHAP uses a game-theoretic framework to attribute feature importance. GuidedBackprop highlights salient input features via positive gradients. Because each technique might emphasise different aspects of the data, benchmarking them was essential for trustworthy interpretation.
@@ III-E2 Remove and Retrain (ROAR)
We systematically compared these methods using the Remove and Retrain (ROAR) framework [15]. First, each interpretability technique ranks the input features by importance. Next, we replace the most influential features (above a chosen threshold) with an uninformative value (zero) [^5] Zero was chosen as the uninformative value because the input features are correlation coefficients and zero correlation indicates that no positive or negative relationship exists between the two variables. Hence, zero is uninformative in the context of functional connectivity.. Finally, we retrain the model on this modified dataset and measure the drop in classification performance. Larger performance decreases indicate that the removed features were indeed critical for the model’s predictive capability.
@@ III-E3 Important Regions of Interest (ROIs)
Figure 2: Loss graphs during first K-Fold. AE1, AE2, and classifier loss represent reconstruction loss during unsupervised pre-training. Model loss and accuracy refer to the supervised fine-tuning stage.
@@
After determining which features (correlation coefficients between brain regions) were most important, we aggregated them at the region level. Specifically, we identified how often each Region of Interest (ROI) appeared among top-weighted features and then mapped these important ROIs to their anatomical labels, visualising the results (Figure 4). We then mapped the features across the different pipelines to Brodmann Areas (Table III) and analysed the overlapping features (Figure 6). This analysis enabled us to pinpoint the key brain areas driving ASD classification and to compare our findings against established ASD-related neurobiology.
@@ IV Results and Discussion
The model achieved state-of-the-art performance on the ABIDE dataset, with a maximum accuracy of 98.2% across three different preprocessing pipelines, exceeding the previous benchmark of 93.59% (Table I). Our overall metrics were equally strong (specificity: 0.98, precision: 0.98, F1-score: 0.97). Below, we detail performance and interpretability results, emphasising the clinical significance of the identified regions.
@@
TABLE I: Accuracy comparison with studies that utilise fMRI modality and have datasets comprising more than 500 samples.
Model Accuracy
Model without FD Filtering 91%
Model with FD Filtering 98.2%
TABLE II: Model accuracy with and without framewise displacement filtering of samples having $>$ 0.2mm FD.
IV-A Model performance
Filtering scans with framewise displacement (FD) >0.2,mm boosted accuracy from 91% to 98.2% (Table II), underscoring the importance of motion correction [51, 50]. This 7% jump aligns with the autoencoders’ struggles to reconstruct high-FD samples, although some studies report minimal effects of FD [52]. Motion correction is thus essential to ensure consistent biomarker detection, an especially relevant point for younger populations where motion artifacts may be more frequent and yet early diagnosis is most critical [8].
Among the four preprocessing pipelines (CCS, CPAC, DPARSF, and NIAK), DPARSF achieved the highest accuracy (98.2%), while CCS and CPAC trailed within 1% and NIAK was about 4% lower (Figure 3). In addition, ABIDE’s multi-center nature introduces site-specific noise from different scanners and protocols. By evaluating four distinct preprocessing pipelines, we partially addressed these variations; however, the relatively lower performance of the NIAK pipeline suggests residual sensitivity to cross-site heterogeneity for NIAK. Future investigations could explore advanced domain adaptation or harmonisation strategies to further mitigate site-based variability.
To check for overfitting, we monitored training and validation loss curves (Figure 2), observing stable convergence without divergence. Moreover, the important ROIs identified by our model closely match independent neuroimaging findings [64, 65], suggesting that it captures genuine ASD-relevant patterns rather than memorising dataset artifacts. Nonetheless, testing on external data is a near-future objective to confirm out-of-distribution generalisation. Observing Figure (2), the autoencoders (AE1 and AE2 loss) consistently reconstruct the data during unsupervised pre-training, suggesting that broad, shared features are captured prior to differentiating ASD from TC. In the subsequent supervised fine-tuning phase, cross-entropy loss (Model loss) drops steadily for both training and validation sets, indicating the network effectively learns the subtle differences between ASD and TC without overfitting. This hybrid approach—unsupervised pre-training plus supervised fine-tuning—appears to capture robust, clinically relevant patterns in functional connectivity, consistent with findings on multi-stage training in neuroimaging tasks [66].
@@ IV-B1 ROAR Benchmark
The ROAR analysis identified which interpretability methods are reliable. It showed that Integrated Gradients, Deep fLift, DeepLiftShap, and GradientShap methods all had similar results and showed the steeper drop in accuracy when the initial features were removed, indicating that the gradient-based methods identified the most discriminative features required for ASD classification (Figure 5). The Integrated Gradients approach showed the steepest drop in accuracy and was therefore assumed the most reliable method. Notably, the LIME approach did not perform much better than the random baseline, indicating that feature perturbation-based techniques do not perform well in the functional connectivity field. This contrasts with results found in the original ROAR study [15] that showed that the Integrated Gradients approach performed similarly to the random baseline, further emphasising that the best-performing interpretability methods change depending on the model architecture and the data’s modality.
@@ IV-B2 Feature Analysis
Feature analysis using the Integrated Gradients approach revealed a set of consistent brain connectivity patterns highly discriminative for the classification of ASD across the different preprocessing pipelines, one of which can be observed in Figure (4).
Important brain areas (parcellated per the AAL atlas), identified by the interpretation method are shown in columns CCS, CPAC, and DPARSF in Table (III). The regions are arranged in descending order of importance score, with their corresponding Brodmann Areas [67]. The most consistently important regions across pipelines are areas 17 (Primary visual cortex) and 18 (Secondary visual cortex), with areas 41 and 42 (Auditory cortex) also showing significance.
@@
The convergence of our machine learning-based findings with these different studies provides strong validation that our model has indeed captured genuine neurobiological markers of ASD, rather than dataset-specific patterns. These results also demonstrate its potential to corroborate and extend existing knowledge in autism research. The consistency of our findings across different preprocessing pipelines and their alignment with diverse research approaches (genetic, neuroanatomical, and functional) highlights the robustness of our method and its potential for advancing our understanding of the neurobiological basis of ASD. Figure (6) highlights these important regions, providing a visual representation of the key areas identified by our analysis.
@@
Figure 6: Shows the Brodmann’s area[72]. The different colors represent different brain functions. The common important brain regions from the Table (III) are highlighted.